# Bridgeland/weak Stability Conditions under Spherical Twist Associated to a Torsion Sheaf

### v1

@@ Abstract.
In this paper, we study the action of an autoequivalence, the spherical twist associated to a torsion sheaf, on the standard Bridgeland stability conditions and a generalized weak stability condition on the derived category of a K3 surface. As a special case, we construct a Bridgeland stability condition associated to a non-nef divisor, which conjecturally lies in the geometric component but outside the geometric chamber. We also discuss the destabilizing objects and stability of certain line bundles at the weak stability condition associated to a nef divisor.
@@ 1. Introduction
Bridgeland stability conditions, introduced by Bridgeland in [3], provide a framework for defining stability conditions in triangulated categories. When the triangulated category is $D^{b}(X)$, the derived category of coherent sheaves on a smooth projective variety, Bridgeland stability conditions have been constructed and extensively studied by many authors (e.g. [4, 1, 2] etc). A Bridgeland stability condition on $D^{b}(X)$ can be equivalently defined as a pair $(Z,\mathcal{A})$, where $\mathcal{A}$ is the heart of a bounded t-structure on $D^{b}(X)$, and $Z$ is a group homomorphism from the Grothendieck group of $\mathcal{A}$ to $\mathbb{H}$. It is shown in [3] that the set of stability conditions forms a complex manifold, on which the autoequivalence group $\mathrm{Aut}(D^{b}(X))$ acts on the left and the universal cover of $\text{GL}^{+}(2,\mathbb{R})$ acts on the right.
Note that in the definition of Bridgeland stability condition, nonzero object in $\mathcal{A}$ can not be mapped to zero by the central charge $Z$. However, as one deforms the central charge towards certain limits, nonzero objects in the heart $\mathcal{A}$ may fall into the kernel of $Z$. This situation gives rise to a notion called weak stability condition, originally defined by [12], and later on used and studied by [2, 5]. These notions allow nonzero objects in $\mathcal{A}$ to be mapped to zero, thus generalizes the definition of Bridgeland stability condition. In [7], the authors further generalized the definition in [12] by allowing more flexibility of the phases of the object in $\ker(Z)$, defining their phases as limits of phases of Bridgeland stability conditions. In this way, the generalized weak stability conditions can be seen as degenerations or “limits” of Bridgeland stability conditions. One can similarly study the group actions on the set of weak stability conditions defined in [7].
The study of group actions on Bridgeland/weak stability conditions has many important applications. For example, in [6], the authors study the action of a specific autoequivalence–the relative Fourier-Mukai transform–on certain Bridgeland/weak stability conditions, and use it to show the stability of line bundles in certain region of the stability manifold. In this application, understanding the image of Bridgeland/weak stability conditions under the action of the relative Fourier-Mukai transform has been demonstrated to be very useful.
@@
Motivated by this application, in this paper, we focus on the action of a different type of autoequivalence, the spherical twist associated with a torsion sheaf. Let $X$ be a K3 surface, we study the action of this spherical twist functor on the standard Bridgeland stability conditions associated to a ample divisor and certain weak stability conditions associated to a nef divisor, and give explicit construction of the images. As a special case of the general construction, we obtain a Bridgeland stability condition associated to a divisor outside the closure of the ample cone, so that the stability condition is outside the geometric chamber, but conjecturally lies in the geometric component of the stability manifold by a conjecture of Bridgeland [4, Conjecture 1.2].
@@ 1.1. Outline
In Section 2, we review the basic definitions and properties of generalized weak stability conditions, including an example associated with a nef divisor, as studied in [7]. We also review the definition and some basics of spherical twist functor on $D^{b}(X)$ for $X$ a K3 surface. In Section 3, we define another weak stability condition associated to a nef divisor. In Section 4, we compute the image of a central charge under the spherical twist functor, and show that the weak stability conditions constructed in Section 3 is the image of the weak stability condition presented in Section 2 under the spherical twist. In Section 5, we construct the image of a standard form Bridgeland stability condition under the spherical twist. Finally, Sections 6 and 7 discuss some results on the stability of line bundles with respect to the weak stability condition in Section 2.
@@ 2.2. Weak stability condition associated to a nef divisor I
In this section we review an example of weak stability constructed in [7]. In 5.13 of [7], we constructed a weak stability condition associated to a particular nef divisor on a Weierstraß elliptic K3 surface, following the construction of Bridgeland stability condition associated to a nef divisor in [15]. The construction in 5.13 of [7] actually works for a more general class of nef divisor on a general K3 surface, so we write it in more generality.
Let $X$ be a K3 surface. Let $C$ be a smooth rational curve such that $C^{2}=-e<0$. Assume that there exist a nef divisor $\nu$ such that $\nu\cdot C=0$ and $\nu\cdot C^{\prime}>0$ for any other irreducible curve $C^{\prime}$ which is not contained in $C$. We consider the central charge $Z_{V,\nu}$ which is defined by
@@
If $B=0$, we omit $B$ in the notation. Recall the construction of heart follows from Tramel-Xia [15]. Set
@@
when $s=0$ we omit $s$ in the notation. We denote the category $\langle\mathcal{O}_{C}(j)|j\leq k\rangle$ by $\mathcal{C}_{k}$, and set
@@
Then [15, Lemma 3.2] implies that $(\mathcal{T}^{1}_{\nu,k},\mathcal{F}^{1}_{\nu,k})$ is a torsion pair in $\mathrm{Coh}^{\nu,s}$, tilting at this torsion pair we obtain the heart
@@
From Proposition 10.6 in [6], we know that there is an equivalent definition of $\mathcal{B}_{\nu,k}$: Proposition 2.1. ([6] Proposition 10.6)  $\mathcal{B}_{\nu,k}$ is obtained by tilting $\mathrm{Coh}(X)$ at the torsion pair
@@
Since $X$ is a K3 surface, we still have the following Proposition in [7]. Proposition 2.2. ([7] Proposition 5.14) We have $Z_{V,\nu}(\mathcal{B}_{\nu,k})\in\mathbb{H}_{0}$ for $k=-1,-2$. Furthermore, we have when $k=-1$,
@@
To define a weak stability condition, we still need to define the phase of the objects in the kernel of the central charge. We only consider the case where $k=-2$. Recall that if $\omega$ is an ample class, then $\nu_{\epsilon}:=\nu+\epsilon\omega$ is ample for $\epsilon>0$. Picking an arbitrary ample class $\omega$, we would like to define weak stability conditions $\sigma^{b}_{V,\nu}=(Z_{V,\nu},\mathcal{B}_{\nu,-2},\{\phi^{b}_{V,\nu}(K)\}_{K\in\ker(Z_{V,\nu})\cap\mathcal{B}_{\nu,-2}})$ as the limit of the standard Bridgeland stability conditions $\sigma_{V,\nu_{\epsilon}}=(Z_{V,\nu_{\epsilon}},\mathrm{Coh}^{\nu_{\epsilon}})$. In particular, we define the phases $\{\phi^{b}_{V,\nu}(K)\}_{K\in\ker(Z_{V,\nu})\cap\mathcal{B}_{\nu,-2}}$ by taking the limit of the phases of $\sigma_{V,\nu_{\epsilon}}(K)$ as $\epsilon\to 0$. Note that
@@
Remark 2.3.  Note that the above definition of the phases of objects in $\ker(Z_{V,\nu})$ as the limits of Bridgeland stability conditions is not unique. E.g. let $B\in\text{NS}_{\mathbb{Q}}(X)$ such that $B\cdot\nu=0$, (e.g. taking $B=uC$). Then [15] shows that the pair $(Z_{V,\nu,B},\mathcal{B}_{\nu,-2})$ defines a Bridgeland stability condition for $-1<C\cdot B<0$ ($0<u<\frac{1}{2}$). We denote this set of Bridgeland stability conditions by $\sigma_{V,\nu,B}$.
@@
Spherical twist functor is defined in [14]. Let $X$ be a projective variety of dimension $n$. Recall that an object $E$ in $D^{b}(X)$ is called spherical if $\textnormal{Ext}^{i}(E,E)\simeq\mathbb{C}$ when $i=0,n$, and equals to 00 otherwise, and $E\otimes\omega_{X}\simeq E$. Definition 2.1. ([14])  Let $E$ be a spherical object in $D^{b}(X)$. The spherical twist functor $\text{ST}_{E}:D^{b}(X)\to D^{b}(X)$ is defined to be the Fourier-Mukai transform with kernel equals to
@@
It is shown in Theorem 1.2 in [14] that the functor $\text{ST}_{E}$ associated to a spherical object $E$ is an exact self-equivalence of $D^{b}(X)$.
@@ 3. Weak stability condition associated to a nef divisor II
By definition, we have $\textnormal{Hom}(\mathcal{T}_{C},\mathcal{F}_{C})=0$. We only need to show for any object $E\in\mathcal{B}_{\nu,-2}$, $E$ fits in a s.e.s.
@@
for some $j\leq-1$. If $H^{-1}(I_{1})=0$, then $H^{0}(I_{1})\neq 0$ and the first s.e.s. in equation 3.1.1 is a s.e.s. in $\mathrm{Coh}(X)$, this implies that $K_{1}\in\mathcal{C}_{-2}$ which is a contradiction. Then $H^{-1}(I_{1})\neq 0$. Then $\mu_{\nu}(K_{1})\geq 0$ and $\mu_{\nu}(H^{-1}(I_{1}))\leq 0$ forces $\mu_{\nu}(K_{1})=0$, hence $K_{1}$ is a torsion sheaf supported on $C$. This implies that $H^{-1}(I_{1})$ is a torsion sheaf supported on $C$ and hence $H^{-1}(I_{1})\in\mathcal{C}_{-2}$. Since $\textnormal{Hom}(K_{1},\mathcal{C}_{-2})=0$ we must have $j=-1$. But GRR implies that $\textnormal{Ext}^{1}(\mathcal{O}_{C}(-1),\mathcal{O}_{C}(-2))=0$, also $\textnormal{Hom}(K_{1},\mathcal{C}_{-2})=0$, hence $\textnormal{Hom}(H^{-1}(I_{1}),\mathcal{O}_{C}(-2))=0$. This contradicts $H^{-1}(I_{1})\in\mathcal{C}_{-2}$.
@@
Let $\mathcal{A}_{\nu,-2}:=\langle\mathcal{F}_{C},\mathcal{T}_{C}[-1]\rangle$. Since $\ker Z_{V,\nu}\cap\mathcal{A}_{\nu,-2}=\langle\mathcal{O}_{C}(-1)[-1]\rangle$, the central charge $Z_{V,\nu}$ also defines a weak stability function on $\mathcal{A}_{\nu,-2}$. We define the phases of the object $\mathcal{O}_{C}(-1)[-1]$ to be $\phi^{a}_{V,\nu}(\mathcal{O}_{C}(-1)[-1])=\frac{1}{2}$. Since $\ker(Z_{V,\nu})\cap\mathcal{A}_{\nu,-2}=\langle\mathcal{O}_{C}(-1)[-1]\rangle$, then any object $K\in\ker(Z_{V,\nu})\cap\mathcal{A}_{\nu,-2}$ has phase equal to $\frac{1}{2}$, and the weak see-saw property is automatically satisfied. Proposition 3.2.The triple $\sigma^{a}_{V,\nu}:=(Z_{V,\nu},\mathcal{A}_{\nu,-2},\{\phi^{a}_{V,\nu}(K)\})$ satisfies the HN property, hence defines a weak stability condition.
@@ 4.1. Central charge formula
As a special case of solving $\omega$ and $B$ from ${\overline{\omega}}$, ${\overline{B}}$, we have the following Proposition. Proposition 4.1. Assume that $\overline{\psi}=0$. Let $g=\begin{pmatrix}1&0\\ 0&\frac{1}{D_{\overline{\omega}}+1}\end{pmatrix}.$
When $t=-1$, we have
@@
When $t=-1$ $D_{\overline{\omega}}=0$, and $\overline{B}=\overline{u}C$.
@@ 4.2. Correspondence of hearts
In this subsection, we check that the correspondence of hearts under $\text{ST}_{\mathcal{O}_{C}(-1)}$. The following result is stated and used in [15], we give the proof here just for completion. Lemma 4.1. Let $\mathcal{C}:=\langle\mathcal{O}_{C}(j)|j\in\mathbb{Z}\rangle$. Assume that $E\in\mathcal{C}$, and $\textnormal{Hom}(\mathcal{O}_{C}(k),E)=0$, then $E\in\mathcal{C}_{k-1}$.
@@
The statement is obviously true when $n=1$. Now we assume that the statement is true for $n\leq l$. Assume that $\mathrm{ch}_{1}(E)=(l+1)C$. Let $i:C\hookrightarrow X$. Note that $i_{*}i^{*}E\simeq F\oplus Q$ where $0\neq F\simeq\oplus_{p}\mathcal{O}_{C}(n_{p})$ and $Q$ is supported in dimension 00. We take $E\to F$ and complete it into a s.e.s. in $\mathrm{Coh}(X)$:
$0\to G\to E\to F\to 0$
@@
for some $q\leq k-1$. Hence if we write $i^{*}G\simeq\oplus_{i}\mathcal{O}_{C}(m_{i})\oplus R$ where $R$ supported in dimension 00, then there exists $i_{0}$ such that $m_{i_{0}}\leq k-1$. Direct computation shows that $L^{-1}i^{*}(F)\simeq\oplus\mathcal{O}_{C}(n_{p}+2)$. Then equation 4.1.1 implies that there exists $p_{0}$ such that $n_{p_{0}}+2\leq k-1$. Take the s.e.s.
@@
We abbreviate notation and denote $\text{ST}_{\mathcal{O}_{C}(-1)}$ by ST. Theorem 4.2. Using the above notation, we have $\text{ST}(\mathcal{B}_{\nu,-2})=\mathcal{A}_{\nu,-2}$.

### v2

@@ Abstract.
In this paper, we study the action of an autoequivalence, the spherical twist associated to a torsion sheaf, on the standard Bridgeland stability conditions and a generalized weak stability condition on the derived category of a K3 surface. As a special case, we construct a Bridgeland stability condition associated to a non-nef divisor, which is outside the geometric chamber, but conjecturally lies in the geometric component of the stability manifold by a conjecture of Bridgeland. We also briefly discuss the stability of certain line bundles at the weak stability condition associated to a nef divisor.
@@ 1. Introduction
Bridgeland stability conditions, introduced by Bridgeland in [4], provide a framework for defining stability conditions in triangulated categories. When the triangulated category is $D^{b}(X)$, the derived category of coherent sheaves on a smooth projective variety, Bridgeland stability conditions associated to an ample divisor have been constructed and extensively studied by many authors (e.g. [5, 1, 3] etc). A Bridgeland stability condition on $D^{b}(X)$ can be equivalently defined as a pair $(Z,\mathcal{A})$, where $\mathcal{A}$ is the heart of a bounded t-structure on $D^{b}(X)$, and $Z$ is a group homomorphism from the Grothendieck group of $\mathcal{A}$ to $\mathbb{H}$. It is shown in [4] that the set of stability conditions forms a complex manifold, on which the autoequivalence group $\mathrm{Aut}(D^{b}(X))$ acts on the left and the universal cover of $\text{GL}^{+}(2,\mathbb{R})$ acts on the right. In the definition of Bridgeland stability condition, nonzero object in $\mathcal{A}$ can not be mapped to zero by the central charge $Z$. However, as one deforms the central charge towards certain limits, nonzero objects in the heart $\mathcal{A}$ may fall into the kernel of $Z$. This situation gives rise to a notion called weak stability condition, originally defined by [14], and later on used and studied by [3, 6]. In [8], the authors further generalized the definition in [14] by allowing more flexibility of the phases of the object in $\ker(Z)$, defining their phases as limits of phases of Bridgeland stability conditions. In this way, the generalized weak stability conditions can be seen as degenerations or “limits” of Bridgeland stability conditions. One can similarly study the group actions on the set of weak stability conditions defined in [8].
@@
The study of group actions on Bridgeland/weak stability conditions has many important applications. For example, in [7], the authors study the action of a specific autoequivalence–the relative Fourier-Mukai transform–on certain Bridgeland/weak stability conditions, and use it to show the stability of line bundles in certain region of the stability manifold. In this application, understanding the image of Bridgeland/weak stability conditions under the action of the relative Fourier-Mukai transform has been demonstrated to be very useful. From the perspective of applications along this line, it is also evident that the negative curves on a surface play a significant role in studying the stability of line bundles, which motivates the exploration of the weak stability condition at the boundary corresponding to negative curves. Motivated by this application, in this paper, we focus on the action of a different type of autoequivalence, the spherical twist associated with a torsion sheaf supported on a negative curve. Let $X$ be a K3 surface, we study the action of this spherical twist functor on the standard Bridgeland stability conditions associated to a ample divisor and certain weak stability conditions associated to a nef divisor, and give explicit construction of the images. As a special case of the general construction, we give the first construction of a Bridgeland stability condition associated to a divisor outside the closure of the ample cone, so that the stability condition is outside the geometric chamber, but conjecturally lies in the geometric component of the stability manifold by a conjecture of Bridgeland [5, Conjecture 1.2]. Finally, we discuss the stability of line bundles with respect to certain weak stability conditions on specific K3 surfaces. Although this is not a direct application of the behavior of stability conditions under spherical twists, understanding the stability of line bundles in weak stability conditions has significant applications to the study of the stability of the line bundles at the standard Bridgeland stability conditions, so we include a brief discussion here.
@@ 1.1. Outline
Since the image of the weak stability condition associated to a nef divisor under spherical twist is easier to describe, we start with the study of weak stability conditions. In Section 2, we review the basic definitions and properties of generalized weak stability conditions, including an example associated with a nef divisor, as studied in [8]. We also review the definition and some basics of spherical twist functor on $D^{b}(X)$ for $X$ a K3 surface. In Section 3, we define another weak stability condition associated to a nef divisor. In Section 4, we compute the image of a central charge under the spherical twist functor, and show that the weak stability conditions constructed in Section 3 is the image of the weak stability condition presented in Section 2 under the spherical twist. In Section 5, we give a construction of the image of a standard form Bridgeland stability condition associated to an ample divisor under the spherical twist. Finally, Sections 6 briefly discuss some results on the stability of line bundles with respect to the weak stability condition in Section 2.
@@ 2.2. Weak stability condition associated to a nef divisor I
In 5.13 of [8], we constructed a weak stability condition associated to a particular nef divisor on a Weierstraß elliptic K3 surface, following the construction of Bridgeland stability condition associated to a nef divisor in [17]. The construction in 5.13 of [8] actually works for a more general class of nef divisor on a general K3 surface, so we write it in more generality.
Let $X$ be a K3 surface with a smooth rational curve $C$ such that $C^{2}=-2$. Assume that there exist a nef divisor $\nu$ such that $\nu\cdot C=0$ and $\nu\cdot C^{\prime}>0$ for any other irreducible curve $C^{\prime}$ which is not contained in $C$. We consider the central charge $Z_{V,\nu}$ which is defined by
@@
If $B=0$, we omit $B$ in the notation. Recall the construction of heart in Tramel-Xia [17]. Set
@@
when $s=0$ we omit $s$ in the notation. Denoting the category $\langle\mathcal{O}_{C}(j)|j\leq k\rangle$ by $\mathcal{C}_{k}$, and set
@@
Then [17, Lemma 3.2] implies that $(\mathcal{T}^{1}_{\nu,k},\mathcal{F}^{1}_{\nu,k})$ is a torsion pair in $\mathrm{Coh}^{\nu,s}$, tilting at this torsion pair we obtain the heart
@@
The category $\mathcal{B}_{\nu,k}$ can be equivalently defined by
Proposition 2.1. ([7] Proposition 10.6)  $\mathcal{B}_{\nu,k}$ is obtained by tilting $\mathrm{Coh}(X)$ at the torsion pair
@@
Since $X$ is a K3 surface, we still have the following Proposition in [8].
Proposition $2.$ ([8] Proposition 5.14) We have $Z_{V,\nu}(\mathcal{B}_{\nu,k})\in\mathbb{H}_{0}$ for k=1,2k=-1,-2. Furthermore, we have when k=1k=-1,
@@
To define a weak stability condition, we still need to define the phase of the objects in the kernel of the central charge. We only consider the case where k=2k=-2. Recall that if $\omega$ is an ample class, then $\nu_{\epsilon}:=\nu+\epsilon\omega$ is ample for $\epsilon>0$. Picking an arbitrary ample class $\omega$, we would like to define weak stability conditions $\sigma^{b}_{V,\nu}=(Z_{V,\nu},\mathcal{B}_{\nu,-2},\{\phi^{b}_{V,\nu}(K)\}_{K\in\ker(Z_{V,\nu})\cap\mathcal{B}_{\nu,-2}})$ as the limit of the standard Bridgeland stability conditions $\sigma_{V,\nu_{\epsilon}}=(Z_{V,\nu_{\epsilon}},\mathrm{Coh}^{\nu_{\epsilon}})$. In particular, we define the phases $\{\phi^{b}_{V,\nu}(K)\}_{K\in\ker(Z_{V,\nu})\cap\mathcal{B}_{\nu,-2}}$ by taking the limit of the phases of $\sigma_{V,\nu_{\epsilon}}(K)$ as $\epsilon\to 0$. Note that
@@
Remark 2.3.  Note that the above definition of the phases of objects in $\ker(Z_{V,\nu})$ as the limits of Bridgeland stability conditions is not unique. E.g. let $B\in\text{NS}_{\mathbb{Q}}(X)$ such that $B\cdot\nu=0$, (e.g. taking $B=uC$). Then [17] shows that the pair $(Z_{V,\nu,B},\mathcal{B}_{\nu,-2})$ defines a Bridgeland stability condition for $1<C\cdot B<0$ ($0<u<\frac{1}{2}$). We denote this set of Bridgeland stability conditions by $\sigma_{V,\nu,B}$.
@@
Spherical twist functor is defined in [16]. Let $X$ be a projective variety of dimension $n$. Recall that an object $E$ in $D^{b}(X)$ is called spherical if $\textnormal{Ext}^{i}(E,E)\simeq\mathbb{C}$ when $i=0,n$, and equals to 0 otherwise, and $E\otimes\omega_{X}\simeq E$.
Definition 2.1. ([16])  Let $E$ be a spherical object in $D^{b}(X)$. The spherical twist functor $\text{ST}_{E}:D^{b}(X)\to D^{b}(X)$ is defined to be the Fourier-Mukai transform with kernel equals to
@@
Theorem 1.2 in [16] implies that the functor $\text{ST}_{E}$ associated to a spherical object $E$ is an exact self-equivalence of $D^{b}(X)$.
@@ 3. Weak stability condition associated to a nef divisor II
By definition, we have $\textnormal{Hom}(\mathcal{T}_{C},\mathcal{F}_{C})=0$. We only need to show for any object $\in\mathcal{B}_{\nu,-2}, EE$ fits in a s.e.s.
@@
for some $j\leq-1$. If $H^{-1}(I_{1})=0$, then $H^{0}(I_{1})\neq 0$ and the first s.e.s. in equation 3.1.1 is a s.e.s. in $\mathrm{Coh}(X)$, this implies that $K_{1}\in\mathcal{C}_{-2}$ which is a contradiction. Then $H^{-1}(I_{1})\neq 0$. Then $\mu_{\nu}(K_{1})\geq 0$ and $\mu_{\nu}(H^{-1}(I_{1}))\leq 0$ forces $\mu_{\nu}(K_{1})=0$, hence $K_{1}$ is a torsion sheaf supported on $C$. This implies that $H^{-1}(I_{1})$ is a torsion sheaf supported on $C$ and hence $H^{-1}(I_{1})\in\mathcal{C}_{-2}$. Since $\textnormal{Hom}(K_{1},\mathcal{C}_{-2})=0$ we must have j=1j=-1. But GRR implies that $\textnormal{Ext}^{1}(\mathcal{O}_{C}(-1),\mathcal{O}_{C}(-2))=0$, also $\textnormal{Hom}(K_{1},\mathcal{C}_{-2})=0$, hence $\textnormal{Hom}(H^{-1}(I_{1}),\mathcal{O}_{C}(-2))=0$. This contradicts $H^{-1}(I_{1})\in\mathcal{C}_{-2}$.
@@
Let $\mathcal{A}_{\nu,-2}:=\langle\mathcal{F}_{C},\mathcal{T}_{C}[-1]\rangle$. Since $\ker Z_{V,\nu}\cap\mathcal{A}_{\nu,-2}=\langle\mathcal{O}_{C}(-1)[-1]\rangle$, the central charge $Z_{V,\nu}$ also defines a weak stability function on $\mathcal{A}_{\nu,-2}$. We define the phases of the object $\mathcal{O}_{C}(-1)[-1]$ to be $\phi^{a}_{V,\nu}(\mathcal{O}_{C}(-1)[-1])=\frac{1}{2}$. Since $\ker(Z_{V,\nu})\cap\mathcal{A}_{\nu,-2}=\langle\mathcal{O}_{C}(-1)[-1]\rangle$, then any object $K\in\ker(Z_{V,\nu})\cap\mathcal{A}_{\nu,-2}$ has phase equal to $\frac{1}{2}$, and the weak see-saw property is automatically satisfied.
Proposition 3.2.The triple $\sigma^{a}_{V,\nu}:=(Z_{V,\nu},\mathcal{A}_{\nu,-2},\{\phi^{a}_{V,\nu}(K)\})$ satisfies the HN property, hence defines a weak stability condition.
@@ 4.1. Central charge formula
As a special case of solving $\omega$ and $B$ from $\overline{\omega}$}, $\overline{B}$}, we have the following Proposition.
Proposition 4.1. Assume that $\overline{\psi}=0$, then $\psi=0$. Let $g=\begin{pmatrix}1&0\\ 0&\frac{1}{D_{\overline{\omega}}+1}\end{pmatrix}$.
When t=1t=-1, we have
@@
When t=1t=-1 $D_{\overline{\omega}}=0$, and $\overline{B}=\overline{u}C$.
@@ 4.2. Correspondence of hearts
In this subsection, we check that the correspondence of hearts under $\text{ST}_{\mathcal{O}_{C}(-1)}$. The following result is stated and used in [17], we give the proof here just for completion.
Lemma 4.1. Let $\mathcal{C}:=\langle\mathcal{O}_{C}(j)|j\in\mathbb{Z}\rangle$. Assume that $E\in\mathcal{C}$, and $\textnormal{Hom}(\mathcal{O}_{C}(k),E)=0$, then $E\in\mathcal{C}_{k-1}$.
@@
The statement is obviously true when $n=1$. Now we assume that the statement is true for $n\leq l$. Assume that $\mathrm{ch}_{1}(E)=(l+1)C$. Let $i:C\hookrightarrow X$. Note that $i_{*}i^{*}E\simeq F\oplus Q$ where $0\neq F\simeq\oplus_{p}\mathcal{O}_{C}(n_{p})$ and $Q$ is supported in dimension 0. We take $E\to F$ and complete it into a s.e.s. in $\mathrm{Coh}(X)$:
$0\to G\to E\to F\to 0.$
@@
for some $q\leq k-1$. Hence if we write $i^{*}G\simeq\oplus_{i}\mathcal{O}_{C}(m_{i})\oplus R$ where $R$ supported in dimension 0, then there exists $i_{0}$ such that $m_{i_{0}}\leq k-1$. Direct computation shows that $L^{-1}i^{*}(F)\simeq\oplus\mathcal{O}_{C}(n_{p}+2)$. Then equation 4.1.1 implies that there exists $p_{0}$ such that $n_{p_{0}}+2\leq k-1$. Take the s.e.s.
@@
We abbreviate notation and denote $\text{ST}_{\mathcal{O}_{C}(-1)}$ by ST.
Theorem 4.2. Using the above notation, we have $\text{ST}(\mathcal{B}_{\nu,-2})=\mathcal{A}_{\nu,-2}$.